# Understanding Impact of Human Feedback via Influence Functions

### v2

@@ Acknowledgments
The authors would like to thank Juyong Lee, Kyuyoung Kim, and Roger Grosse for helpful comments on the project. We thank Hanho Ryu for providing helpful baseline experiments regarding Mahalanobis and KNN. We thank Seungku Kim and Sinjae Kang for providing human labels for our human surveys. This work was supported by National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00256259), Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)) and Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2025-02304967, AI Star Fellowship(KAIST)). We also thank the support from Google Cloud through the Google Gemma 2 Academic Program GCP Credit Award.
@@ A.1 Vector Compression Method
OPORP is a straightforward two-step method consisting of (1) permutation and (2) projection. In the first step, the gradient vector is permuted using a permutation matrix. Specifically, we implement the efficient permutation technique proposed in  Lin et al. (2024), where the vector is permuted using multiple sub-permutations. In the second step, the permuted gradient vector undergoes element-wise multiplication with a projection vector, denoted as $\rho$, where each element $\rho_{i}$ is randomly sampled from ${-1,+1}$ with equal probability.
@@ B.1 Bias Detection
Table 4: Details on datasets used in bias detection. The “Data split” indicates the dataset split of Anthropic-HH.
@@ B.2 Labeling Strategy Oversight
Answer: {tblr} colspec = @X@, rowsep=1pt, System prompt: You are a careful, helpful and diligent assistant. Your task is to evaluate conversations between a human and an AI assistant, and you will evaluate which of two responses is better in terms of helpfulness.
@@ C.1 LLM Baselines
In the length bias experiment, the 3-shot examples consisted of conversations where the chosen response was deliberately more concise, depending on the prompt structure. The model was tasked with evaluating both the brevity and the quality of the content. Additionally, in the sycophancy bias experiment, the 3-shot examples included scenarios where the chosen response was factually accurate but less aligned with the user’s opinion, while the rejected response exhibited sycophantic tendencies. The prompts used for 3-shot learning in these LLM experiments can be found in  subsection B.2, and the few shot examples used can be found in  Figure 6 and  Figure 7
@@ C.2.1 Mahalanobis Distance
Mahalanobis $\text{Mahalanobis distance}=d_{M}(X_{\text{act}})$
@@ 3. The output should be a JSON object of the following format:
{"$\text{feedback}":<\text{write a feedback for criteria}>,"\text{score}":<\text{an integer number between 1 and 5}>\}$
@@
Output: {tblr} colspec = @X@, rowsep=1pt, Task Description: An instruction, a response to evaluate, and a score rubric representing an evaluation criteria are given.
@@
We adopt the direct assessment prompt of Prometheus2 (Kim et al., 2024) to construct our sycophancy score labeling prompt. Our prompt queries Gemini-1.5-pro to rate a Likert scale score ranging from 1 to 5 regarding a scoring rubric that gives a detailed explanation of how to rate sycophancy scores for responses. We have tested various wordings or phrases and selected the prompt with the highest correlation with human evaluation. We provide the resulting prompt in Appendix D.
@@ Appendix F Alice and Bob Experiment Weight Update Method
     $\mathbf{r}_{i}\leftarrow\mathbf{r}^{(z_{i})}-\mathbf{r}^{(1-z_{i})}$triangleright Subtract scores of losing from winning
     $\eta\leftarrow\text{median of }\mathcal{I}_{\text{val}}(\mathbf{d}_{i})$triangleright Ensure 50:50 ratio of $t_{i}=0$ and $t_{i}=1$
     $t_{i}\leftarrow\mathbb{I}\left[\mathcal{I}_{\text{val}}(\mathbf{d}_{i})\leq\eta\right]$triangleright Large influence considered as mislabeling
@@ Appendix G Reward Retraining on Curated Set Using Bias Detection
In this section, we evaluate whether flipping the preference labels of negatively-contributing samples detected by influence functions improves reward model performance, measured by validation set accuracy. We compare retrained models against reward models trained on the datasets from our main analysis (referred to as main datasets). Our main datasets contain manually flipped preference labels. We also compare them with a clean dataset without any label modifications. Specifically, we flip $\alpha$-($\alpha={50,100,150}$) of the actual number of manually flipped preference labels from the training dataset. For example, for $\alpha=50$% we flip the top 3.28% samples with the highest influence scores as the length training dataset contains 6.56% of manually flipped preference labels. We retrain the reward model using this curated dataset and then report its validation set accuracy. For length bias experiments, Concise and Verbose validation set accuracies are reported. For sycophancy bias experiments, Less Sycophantic and More Sycophantic validation set accuracies are reported. Please refer to Appendix B for details on these validation sets.
@@
+ Flipped ($3.28$%) 54.54 (+4.61) 42.84 (-2.09)
+ Flipped ($6.56$%) 59.37 (+9.44) 40.78 (-4.15)
+ Flipped ($9.84$%) 64.12 (+14.20) 38.07 (-6.86)
@@
+ Flipped ($2.09$%) 61.99% (+6.43) 73.33 (-0.67)
+ Flipped ($4.17$%) 61.99% (+6.43) 69.33 (-4.67)
+ Flipped ($6.26$%) 71.93% (+16.37) 68.00 (-6.00)
@@
Table 10: Comparison of reward model validation accuracy after curating the training dataset. "Flipped ($r$%)" indicates that $r$of the dataset were flipped. Length (main) / Sycophancy (main) indicates the datasets with manually injected bias used in our main analysis of bias detection. Clean dataset indicates a dataset without manually injected bias. The symbol ($\pm$%) indicates the difference to Length Original and Sycophancy Original.
@@ I.4 Validation Distribution Ablation using HelpSteer2 Validation Set
Anthropic-HH Anthropic-HH 0.800
Anthropic-HH HelpSteer2 0.620

### v3

@@ Acknowledgments
The authors would like to thank Juyong Lee, Kyuyoung Kim, and Roger Grosse for helpful comments on the project. We thank Hanho Ryu for providing helpful baseline experiments regarding Mahalanobis and KNN. We thank Seungku Kim and Sinjae Kang for providing human labels for our human surveys. This work was supported by National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00414822), Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)) and Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2025-02304967, AI Star Fellowship(KAIST)). We also thank the support from Google Cloud through the Google Gemma 2 Academic Program GCP Credit Award.
@@ A.1 Vector Compression Method
OPORP is a straightforward two-step method consisting of (1) permutation and (2) projection. In the first step, the gradient vector is permuted using a permutation matrix. Specifically, we implement the efficient permutation technique proposed in  Lin et al. (2024), where the vector is permuted using multiple sub-permutations. In the second step, the permuted gradient vector undergoes element-wise multiplication with a projection vector, denoted as $\rho$, where each element $\rho_{i}$ is randomly sampled from 1,+1{-1,+1} with equal probability.
@@ B.1 Bias Detection
Table 4: Details on datasets used in bias detection. The “Data split” indicates the dataset split of Anthropic-$H$.
@@ B.2 Labeling Strategy Oversight
System prompt: You are a careful, helpful and diligent assistant. Your task is to evaluate conversations between a human and an AI assistant, and you will evaluate which of two responses is better in terms of helpfulness.
@@
Answer:
@@ C.1 LLM Baselines
In the length bias experiment, the 3-shot examples consisted of conversations where the chosen response was deliberately more concise, depending on the prompt structure. The model was tasked with evaluating both the brevity and the quality of the content. Additionally, in the sycophancy bias experiment, the 3-shot examples included scenarios where the chosen response was factually accurate but less aligned with the user’s opinion, while the rejected response exhibited sycophantic tendencies. The prompts used for 3-shot learning in these LLM experiments can be found in  Table 6, and the few shot examples used can be found in  Figure 6 and  Figure 7
@@ C.2.1 Mahalanobis Distance
$\text{Mahalanobis distance}=d_{M}(X_{\text{act}})$
@@ Appendix D Sycophancy Bias Labeling Prompt and Details
Task Description: An instruction, a response to evaluate, and a score rubric representing an evaluation criteria are given.
@@ 3. The output should be a JSON object of the following format:
[…]
@@
Output:
@@
We adopt the direct assessment prompt of Prometheus2 (Kim et al., 2024) to construct our sycophancy score labeling prompt. Our prompt queries Gemini-1.5-pro to rate a Likert scale score ranging from 1 to 5 regarding a scoring rubric that gives a detailed explanation of how to rate sycophancy scores for responses. We have tested various wordings or phrases and selected the prompt with the highest correlation with human evaluation. We provide the resulting prompt in Table 8.
@@ Appendix F Alice and Bob Experiment Weight Update Method
  $\mathbf{r}_{i}\leftarrow\mathbf{r}^{(z_{i})}-\mathbf{r}^{(1-z_{i})}$ $\triangleright$ Subtract scores of losing from winning
  $\eta\leftarrow\text{median of }\mathcal{I}_{\text{val}}(\mathbf{d}_{i})$ $\triangleright$ Ensure 50:50 ratio of $t_{i}=0$ and $t_{i}=1$
  $t_{i}\leftarrow\mathbb{I}\left[\mathcal{I}_{\text{val}}(\mathbf{d}_{i})\leq\eta\right]$ $\triangleright$ Large influence considered as mislabeling
@@ Appendix G Reward Retraining on Curated Set Using Bias Detection
In this section, we evaluate whether flipping the preference labels of negatively-contributing samples detected by influence functions improves reward model performance, measured by validation set accuracy. We compare retrained models against reward models trained on the datasets from our main analysis (referred to as main datasets). Our main datasets contain manually flipped preference labels. We also compare them with a clean dataset without any label modifications. Specifically, we flip $\alpha$-% ($\alpha={50,100,150}$) of the actual number of manually flipped preference labels from the training dataset. For example, for $\alpha=50$% we flip the top 3.28% samples with the highest influence scores as the length training dataset contains 6.56% of manually flipped preference labels. We retrain the reward model using this curated dataset and then report its validation set accuracy. For length bias experiments, Concise and Verbose validation set accuracies are reported. For sycophancy bias experiments, Less Sycophantic and More Sycophantic validation set accuracies are reported. Please refer to Appendix B for details on these validation sets.
@@
+ Flipped (3.283.28%) 54.54 (+4.61) 42.84 (-2.09)
+ Flipped (6.566.56%) 59.37 (+9.44) 40.78 (-4.15)
+ Flipped (9.849.84%) 64.12 (+14.20) 38.07 (-6.86)
@@
+ Flipped (2.092.09%) 61.99% (+6.43) 73.33 (-0.67)
+ Flipped (4.174.17%) 61.99% (+6.43) 69.33 (-4.67)
+ Flipped (6.266.26%) 71.93% (+16.37) 68.00 (-6.00)
@@
Table 10: Comparison of reward model validation accuracy after curating the training dataset. "Flipped (rr%)" indicates that rr% of the dataset were flipped. Length (main) / Sycophancy (main) indicates the datasets with manually injected bias used in our main analysis of bias detection. Clean dataset indicates a dataset without manually injected bias. The symbol ($\pm$%) indicates the difference to Length Original and Sycophancy Original.
@@ I.4 Validation Distribution Ablation using HelpSteer2 Validation Set
Anthropic-$H$ Anthropic-$H$ 0.800
Anthropic-$H$ HelpSteer2 0.620